# Sequences with Inequalities

### v1

@@ Abstract.
We consider infinite sequences of positive numbers. The connection between log-concavity and the Bessenrodt–Ono inequality had been in the focus of several papers. This has applications in the white noise distribution theory and combinatorics. We improve a recent result of Benfield and Roy and show that for the sequence of partition numbers $\{p(n)\}$ Nicolas’ log-concavity result implies the result of Bessenrodt and Ono towards $p(n)\,p(m)>p(n+m)$.
@@ 1. Introduction
Nevertheless, non-real rooted polynomials as the chromatic polynomials related to the four-color conjecture have this propery [Hu12] and play an important role in algebraic geometry and Hodge theory to combinatorics [Ka22]. Infinite sequences seem to be not accessable by these methods, although, for example $p(n)$ the number of partitions [On04] of $n$, have also turned out to be log-concave [Ni78, DP15] for all even positive $n$ and all odd numbers $n>25$ proven by the circle method of Hardy–Ramanujan and an exact formula by Rademacher. It has been discovered that (1) is also related to the hyperbolicity of the Jensen polynomials associated with the reciprocal of the Dedekind eta function (the generating function of the $p(n)$) of degree two, in the context of studying the Riemann hypothesis [GORZ19]. Recently, another property had been discovered by Bessenrodt and Ono [BO16], which is a mixture of additive and multiplicative behaviour:
@@
This inequality appears at a first glance as a very special property. But, further sequences turned out to satisfy this Bessenrodt–Ono inequality and similiar inequalities appear as conditions for a white noise distribution theory [AKK00, AKK01]. Initiated by the work by Asai, Kubo, and Kuo there is evidence that under certain conditions log-concavity does imply the Bessenrodt–Ono inequality (BO-inequality), see Corollary 3.
@@
We also refer to ([GMU24], Theorem 4.4). Since many important sequences, as $p(n)$, are not log-concave for all $n$, Benfield and Roy [BR24] came up with an interesting result, motivated by the sequence of partition numbers.
@@
Benfield and Roy applied their result to the sequence $\{p(n)\}_{n=0}^{\infty}$ Since for $N=26$ and $k=0$ the condition (2) is fulfilled, they obtain from the log-concavity property that the BO-inequality holds for $m,n>25$.
Unfortunately, this does not cover the complete result by Bessenrodt–Ono. There are still infinitely many pairs $(m,n)$ left to be checked. For example the BO-inequality holds true for all pairs $(2,n)$, where $n>25$.
@@
[…]
@@
Let the sequence $\{p(n)\}$ of partition numbers be given. Then the log-concavity result by Nicolas implies the BO-inequality result by Bessenrodt and Ono.
Essentially Theorem 1 is still true if we assume
(4) $\alpha(n_{0})^{1/n_{0}}\geq\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}$
and put
$A=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}\geq\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}.$
If we suppose $m,n\geq 1$ satisfy
@@ 2. Proof of Theorem 1
Let $\alpha\left(n\right)$ be the elements of a sequence as in Theorem A. For such a sequence the authors of [AKK00] arrive for $n\geq 0$ and $m\geq 1$ in the first line on page 84 at the inequality $\alpha\left(n\right)\alpha\left(m\right)\leq\alpha\left(0\right)\alpha\left(n+m\right)$. If we now assume $\alpha\left(0\right)<1$ (instead of $\alpha\left(0\right)=1$) this shows
@@ 3. Applications and variants
The following first example shows that our conditions $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left(a\right)\right)^{1/a}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ is sufficient but not necessary.
@@
Let $n_{0}=4$, $\alpha\left({0}\right)$ and $\alpha\left({1}\right)$ positive but arbitrary, $\alpha\left({2}\right)=7$, $\alpha\left({3}\right)=5$, $\alpha\left({4}\right)=15$, and $\alpha\left({n}\right)=\frac{30}{\left(n-4\right)!}$ for $n\geq 5$. We have $\frac{\left(\alpha\left({3}\right)\right)^{2}}{\alpha\left({2}\right)\alpha\left({4}\right)}=\frac{25}{105}<1$, $\frac{\left(\alpha\left({4}\right)\right)^{2}}{\alpha\left({3}\right)\alpha\left({5}\right)}=\frac{225}{150}>1$, $\frac{\left(\alpha\left(5\right)\right)^{2}}{\alpha\left(4\right)\alpha\left(6\right)}=\frac{900}{225}>1$, and $\frac{\left(\alpha\left({n}\right)\right)^{2}}{\alpha\left({n-1}\right)\alpha\left({n+1}\right)}=\frac{n-3}{n-4}>1$ for $n\geq 6$. The sequence consisting of the $\alpha\left({n}\right)$ is log-concave for $n\geq n_{0}=4$. Furthermore, $\left(\alpha\left(2\right)\right)^{2}=49>15=\alpha\left(4\right)$, $\alpha\left({2}\right)\alpha\left({3}\right)=35>30=\alpha\left({5}\right)$, $\alpha\left({2}\right)\alpha\left({4}\right)=105>15=\alpha\left({6}\right)$, $\alpha\left({2}\right)\alpha\left({b}\right)=\frac{210}{\left(b-4\right)!}>\frac{30}{\left(b-2\right)!}=\alpha\left({2+b}\right)$ for $b\geq 5$, $\left(\alpha\left(3\right)\right)^{2}=25>15=\alpha\left(6\right)$, $\alpha\left(3\right)\alpha\left(4\right)=75>5=\alpha\left(7\right)$, $\alpha\left(3\right)\alpha\left(b\right)=\frac{150}{\left(b-4\right)!}>\frac{30}{\left(b-1\right)!}=\alpha\left(3+b\right)$ for $b\geq 5$, $\left(\alpha\left(4\right)\right)^{2}=225>\frac{5}{4}=\alpha\left(8\right)$, $\alpha\left(4\right)\alpha\left(b\right)=\frac{450}{\left(b-4\right)!}>\frac{30}{b!}=\alpha\left(4+b\right)$ for $b\geq 5$, and $\alpha\left(a\right)\alpha\left(b\right)=\frac{900}{\left(a-4\right)!\left(b-4\right)!}>\frac{30}{\left(a+b-4\right)!}=\alpha\left(a+b\right)$ for $a,b\geq 5$. But $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}=15^{1/4}<3=\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left({2}\right)\right)^{1/2}=\sqrt{7}<3=\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}$.
@@
Let $\alpha\left({n}\right)$ be the positive elements of a sequence that is log-concave for all $n\geq 1$ and which satisfies the Bessenrodt–Ono inequality (5)
@@
Now we consider the plane partition numbers $\mathop{\rm pp}\left(n\right)$. By [OPR22] this sequence is log-concave for $n\geq n_{0}=12$. We can check that $\left(\mathop{\rm pp}\left(12\right)\right)^{1/12}>\frac{1479}{859}$. We obtain $\left\{1\leq a\leq 10:\left(\mathop{\rm pp}\left(a\right)\right)^{1/a}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 10\right\}$. Computationally checking the cases for $2\leq a,b\leq 10$ then yields our previous result with Tröger, see [HNT23].
@@
Let $\sum_{n=0}^{\infty}\bar{p}\left(n\right)q^{n}=\prod_{n=1}^{\infty}\frac{1+q^{n}}{1-q^{n}}$ define the overpartitions. By the result of Engel [En17] it is strictly log-concave $\left(\bar{p}\left(n\right)\right)^{2}>\bar{p}\left(n-1\right)\bar{p}\left(n+1\right)$ for $n\geq 3$. Nevertheless it is already log-concave (1) for all $n\geq 1$. We use $n_{0}=4$. Then $\left(\bar{p}\left(n_{0}\right)\right)^{1/n_{0}}>\frac{7}{4}$ and $A=\left\{1\leq a\leq 2:\left(\bar{p}\left(a\right)\right)^{1/a}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}\right\}=\left\{1,2\right\}$. By our result we obtain that the BO-inequality (6) is fulfilled for all $n\geq 3$ or $m\geq 3$. By computationally checking the cases $1\leq n,m\leq 2$ we obtain the result of [LZ21]: the overpartitions fulfill the BO-inequality (6) for all $n,m\geq 1$ except for $\left(n,m\right)\in\left\{\left(1,1\right),\left(1,2\right),\left(2,1\right)\right\}$ where we have equality. Equality also holds when $n=0$ or $m=0$.
@@
    Let $k=2$. By a result by Dong and Ji [DJ24] this is log-concave for $n\geq n_{0}=33$. We obtain $\left(p_{2}\left(33\right)\right)^{1/33}>\frac{224}{195}$ and $A=\left\{1\leq a\leq 31:\left(p_{2}\left(a\right)\right)^{1/a}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}\right\}=\left\{a\in\mathbb{N}_{0}:3\leq a\leq 31\right\}$. By computationally checking the cases $3\leq n,m\leq 31$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the BO-inequality.
@@
    Now let $k=3$. Then again by [DJ24] this is log-concave for $n\geq n_{0}=58$. We obtain $\left(p_{3}\left(58\right)\right)^{1/58}>\frac{525}{463}$ and $A=\left\{1\leq a\leq 56:\left(p_{3}\left(a\right)\right)^{1/a}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 56\right\}$. Computationally checking the cases $2\leq n,m\leq 56$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the BO-inequality.

### v2

@@ Abstract.
We consider infinite sequences of positive numbers. The connection between log-concavity and the Bessenrodt–Ono inequality had been in the focus of several papers. This has applications in the white noise distribution theory and combinatorics. We improve a recent result by Benfield and Roy and show that for the sequence of partition numbers $\{p(n)\}$ Nicolas’ log-concavity result implies the result by Bessenrodt and Ono towards p(n)p(m)>p(n+m)p(n)\,p(m)>p(n+m). We provide several examples. Benfield and Roy gave a conjecture related to $\ell$-ary partition numbers. We prove part of this conjecture.
@@ 1. Introduction
Nevertheless, non-real rooted polynomials as the chromatic polynomials related to the four-color conjecture have this propery [Hu12] and play an important role in algebraic geometry and Hodge theory to combinatorics [Ka22]. Infinite sequences seem to be not accessable by these methods, although, for example $p(n)$ the number of partitions [On04] of $n$, have also turned out to be log-concave [Ni78, DP15] for all even positive $n$ and all odd numbers $n>25$ proven by the circle method by Hardy–Ramanujan and an exact formula by Rademacher. It has been discovered that (1) is also related to the hyperbolicity of the Jensen polynomials associated with the reciprocal of the Dedekind eta function (the generating function of the $p(n)$) of degree two, in the context of studying the Riemann hypothesis [GORZ19]. Recently, another property had been discovered by Bessenrodt and Ono [BO16], which is a mixture of additive and multiplicative behaviour:
@@
This inequality appears at a first glance as a very special property. However, additional sequences were found to satisfy this Bessenrodt–Ono inequality and similiar inequalities appear as conditions for a white noise distribution theory [AKK00, AKK01]. Initiated by the work by Asai, Kubo, and Kuo there is evidence that under certain conditions log-concavity does imply the Bessenrodt–Ono inequality, see Corollary 6.
@@
We also refer to ([GMU25], Theorem 4.4). Since many important sequences, as $p(n)$, are not log-concave for all $n$, Benfield and Roy [BR24a] came up with an interesting result, motivated by the sequence of partition numbers.
@@
Benfield and Roy applied their result to the sequence $\{p(n)\}_{n=0}^{\infty}$. Since for $N=26$ and $k=0$ the condition (2) is fulfilled, they obtain from the log-concavity property that the Bessenrodt–Ono inequality holds for $m,n>25$.
Unfortunately, this did not cover the complete result by Bessenrodt–Ono. There are still infinitely many pairs $(m,n)$ left to be checked. For example the Bessenrodt–Ono inequality holds true for all pairs $(2,n)$, where $n>25$.
@@
Remark 2.
  1. (1)
    The index $n_{0}$ is not unique and may be chosen suitably.
  2. (2)
    After Theorem 1 was published on arXiv, see [HN24], an improved version [BR24b] of Benfield and Roy’s result appeared also covered by our result.
We illustrate the previous theorem with two examples.
Example 3 (see ([BR24b], 1.1, ex. 4)).
Let $\{\alpha(n)\}$ be given by $\alpha\left(0\right)=1/2$ and $\alpha\left(n\right)=2^{\sqrt{n}}$. Then the sequence is log-concave for all $n\geq 1$. Let $n_{0}=2$. Then $\alpha(2)^{1/2}>\alpha(2)/\alpha(1)$. Therefore, condition (3) is fulfilled and we obtain the strict Bessenrodt–Ono inequality as stated in Theorem 1 for $n,m\geq n_{0}-1=1$. In this case, the set AA is empty and plays no role.
Example 4 (see [BR24b], 1.1, ex. 5)).
Define the sequence $\alpha\left(n\right)=\left(1.24\right)^{n}+0.000001$ for $0\leq n\leq 24$ and $\alpha\left(n\right)=p\left(n\right)$ for $n\geq 25$. This sequence is log-concave for $n\geq 26$. Let $n_{0}=27$. Then condition (3) is fulfilled and $A=\left\{2,3,\ldots,25\right\}$. This leads to the strict Bessenrodt–Ono inequality for all $n,m\geq 2$ with $n\geq 26$ or $m\geq n_{0}-1=26$.
Note that taking $n_{0}=26$ we obtain the strict Bessenrodt–Ono inequality only for all $n,m\geq 25$.
In the case of partition numbers $p(n)$ Theorem 1 leads to the following interesting application.
@@
Let the sequence $\{p(n)\}$ of partition numbers be given. Then the log-concavity result by Nicolas [Ni78] implies the Bessenrodt–Ono inequality result by Bessenrodt and Ono [BO16].
Proof.
Let $\alpha(n)=p(n)$ and $n_{0}=26$. Then (3) is satisfied. Moreover,
$A=\{a\in\mathbb{N}_{0}:2\leq a<25\}.$
Therefore, only finitely many pairs given by $\{\left(m,n\right)\in\mathbb{N}_{0}^{2}\,:\,2\leq m,n<25\}$ remain to be checked by a computer calculation. ∎
The sequence $\{c^{n}\}$ for $c>0$ is log-concave but does not satisfy the strict Bessenrodt–Ono inequality. Nevertheless, a minor modification of Theorem 1 can be applied. Let
$\alpha(n_{0})^{1/n_{0}}\geq\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}$
and
$A:=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}\geq\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}.$
If $m,n\geq 1$ satisfy
@@
In the last section we invest into a recent conjecture by Benfield and Roy related to $\ell$-ary partition numbers [BR24b].
@@ 2. Proof of Theorem 1
Let $\alpha\left(n\right)$ be the members of a sequence as in Theorem A. For such a sequence the authors of [AKK00] arrive for $n\geq 0$ and $m\geq 1$ in the first line on page 84 at the inequality $\alpha\left(n\right)\alpha\left(m\right)\leq\alpha\left(0\right)\alpha\left(n+m\right)$. If we now assume $\alpha\left(0\right)<1$ (instead of $\alpha\left(0\right)=1$) this shows
@@ 3. Applications and variants
The following example shows that our conditions
$\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}\text{ and }\left(\alpha\left(a\right)\right)^{1/a}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$
are sufficient but not necessary to obtain the Bessenrodt–Ono inequality.
@@
We consider the sequence given in Table 1 with $\alpha\left({0}\right)$ and $\alpha\left({1}\right)$ positive.
[…]
Table 1. Definition of the sequence from Example 7
The sequence consisting of the $\alpha\left({n}\right)$ is log-concave for $n\geq n_{0}=4$, see Table 2.
[…]
Table 2. Log-concavity and -convexity in Example 7
Furthermore, we obtain the Bessenrodt–Ono inequality for all instances $a,b\geq 2$, see Table 3.
[…]
Table 3. Instances of the Bessenrodt–Ono inequality $\alpha\left(a\right)\alpha\left(b\right)>\alpha\left(a+b\right)$ for Example 7
But for $n_{0}=4$ we obtain $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}=15^{1/4}<3=\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left({2}\right)\right)^{1/2}=\sqrt{7}<3=\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}$.
@@
Let $\alpha\left({n}\right)$ be a sequence of positive numbers that is log-concave for all $n\geq 1$ and which satisfies the Bessenrodt–Ono inequality (4)
@@
Now we consider the plane partition numbers $\mathop{\rm pp}\left(n\right)$. By [OPR22] this sequence is log-concave for $n\geq n_{0}=12$. We can check that $\left(\mathop{\rm pp}\left(12\right)\right)^{1/12}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}$. We obtain
$A=\left\{1\leq a\leq 10:\left(\mathop{\rm pp}\left(a\right)\right)^{1/a}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 10\right\}.$
Computationally checking the cases for $2\leq m,n\leq 10$ then yields our previous result with Tröger, see [HNT23].
@@
Let $\sum_{n=0}^{\infty}\bar{p}\left(n\right)q^{n}=\prod_{n=1}^{\infty}\frac{1+q^{n}}{1-q^{n}}$ define the overpartitions. By the result by Engel [En17] it is strictly log-concave $\left(\bar{p}\left(n\right)\right)^{2}>\bar{p}\left(n-1\right)\bar{p}\left(n+1\right)$ for $n\geq 3$. Nevertheless it is already log-concave (1) for all $n\geq 1$. Let $n_{0}=4$. Then $\left(\bar{p}\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}$ and
$A=\left\{1\leq a\leq 2:\left(\bar{p}\left(a\right)\right)^{1/a}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}\right\}=\left\{1,2\right\}.$
By our result we obtain that the Bessenrodt–Ono inequality (5) is fulfilled for all $n\geq 3$ or $m\geq 3$. By computationally checking the cases $1\leq n,m\leq 2$ we obtain the result of [LZ21]: the overpartitions fulfill the Bessenrodt–Ono inequality (5) for all $n,m\geq 1$ except for $\left(n,m\right)\in\left\{\left(1,1\right),\left(1,2\right),\left(2,1\right)\right\}$ where we have equality. Equality also holds when $n=0$ or $m=0$.
@@
    Let $k=2$. By a result by Dong and Ji [DJ24] this is log-concave for $n\geq n_{0}=33$. We obtain $\left(p_{2}\left(33\right)\right)^{1/33}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}$ and
    $A=\left\{1\leq a\leq 31:\left(p_{2}\left(a\right)\right)^{1/a}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}\right\}=\left\{a\in\mathbb{N}_{0}:3\leq a\leq 31\right\}.$
    By computationally checking the cases $3\leq n,m\leq 31$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the Bessenrodt–Ono inequality.
@@
    Now let $k=3$. Then again by [DJ24] this is log-concave for $n\geq n_{0}=58$. We obtain $\left(p_{3}\left(58\right)\right)^{1/58}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}$ and
    $A=\left\{1\leq a\leq 56:\left(p_{3}\left(a\right)\right)^{1/a}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 56\right\}.$
    Computationally checking the cases $2\leq n,m\leq 56$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the Bessenrodt–Ono inequality.
Proposition 12.
Let $\left\{\alpha\left(n\right)\right\}_{n=0}^{\infty}$ be a sequence of positive numbers such that
$\left(\alpha\left(n\right)\right)^{\frac{1}{n}}>\left(\alpha\left(n+1\right)\right)^{\frac{1}{n+1}}$
for all $n\geq n_{0}\geq 1$. Then $\alpha\left(n\right)\alpha\left(m\right)>\alpha\left(n+m\right)$ for all $n,m\geq n_{0}$.
Proof.
We obtain via a direct computation
$\alpha\left(n\right)\alpha\left(m\right)$ $=$ $\left(\left(\alpha\left(n\right)\right)^{\frac{1}{n}}\right)^{n}\left(\left(\alpha\left(m\right)\right)^{\frac{1}{m}}\right)^{m}$
$>$ $\left(\left(\alpha\left(n+m\right)\right)^{\frac{1}{n+m}}\right)^{n}\left(\left(\alpha\left(n+m\right)\right)^{\frac{1}{n+m}}\right)^{m}=\alpha\left(n+m\right).$
4. On a conjecture by Benfield and Roy
Benfield and Roy made a conjecture ([BR24b], Conjecture 5.1) on the strict Bessenrodt–Ono inequality of $\ell$-ary partition numbers $b^{\ell}\left(n\right)$ of $n$. It is well known ([GMU25], Example 4.7) that the $\ell$-ary sequences are not log-concave. We first invest in the case $\ell=5$. Obviously $b^{5}\left(n\right)b^{5}\left(m\right)\leq b^{5}\left(n+m\right)$ if $m$ or $n\in\left\{1,2,3,4\right\}$. This is similar to case $m$ or $n\in\left\{1\right\}$ in [BO16]. For the other cases they conjectured that $7\leq m\leq 9$ and $6\leq n\leq 9$. But since $b^{5}\left(6\right)=2=b^{5}\left(7\right)$ and $b^{5}\left(13\right)=3$ this has to be modified. The modified conjecture is recorded in Table 4.
[…]
Table 4. Conjectured equality and failure of $b^{5}\left(n\right)b^{5}\left(m\right)\geq b^{5}\left(n+m\right)$ for $n\geq m\geq 1$
Theorem 13.
The conjecture holds true for all pairs $m,n\leq 7500$.
This has been verified with PARI/GP.
The same applies to the $\ell$-ary partition numbers $b^{\ell}\left(n\right)$ (Table 5).
$b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n+m\right)$ $b^{\ell}\left(n\right)b^{\ell}\left(m\right)<b^{\ell}\left(n+m\right)$
$1\leq m\leq\ell-1$ with $1\leq m\leq\ell-1$ with
$n\equiv 0,1,\ldots,\ell-m-1\pmod{\ell}$ $n\equiv\ell-m,\ldots,\ell-1\pmod{\ell}$
$m,n\leq 2\ell-1,m+n\geq 3\ell$
Table 5. Conjectured equality and failure of $b^{\ell}\left(n\right)b^{\ell}\left(m\right)\geq b^{\ell}\left(n+m\right)$ for $n\geq m\geq 1$ and $\ell\geq 5$
In the following we prove that all these cases of equality and failure in Table 5 occur. For $1\leq s\leq\ell$ and $\left(s-1\right)\ell\leq n\leq s\ell-1$ holds $b^{\ell}\left(n\right)=s$. For $0\leq r\leq\ell-1$ also $b^{\ell}\left(s\ell+r\right)=b^{\ell}\left(s\ell\right)$ holds true for all $s\geq 0$ since the smallest part in a partition different from $1$ is $\ell$. For the small values $1\leq m\leq\ell-1$ we obtain therefore $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n\right)=b^{\ell}\left(n+m\right)$ if $n\equiv 0,1,\ldots,\ell-m-1\pmod{\ell}$. On the other hand $b^{\ell}\left(\ell s-1\right)<b^{\ell}\left(\ell s\right)$. Therefore, $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n\right)<b^{\ell}\left(n+m\right)$ if $n\equiv\ell-m,\ell-m+1,\ldots,\ell-1\pmod{\ell}$. Furthermore for $\ell\leq n,m\leq 2\ell-1$ we obtain $2\ell\leq n+m\leq 4\ell-2$ and therefore $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=4>3=b^{\ell}\left(n+m\right)$ if $n+m\leq 3\ell-1$ and $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=4=b^{\ell}\left(n+m\right)$ if $n+m\geq 3\ell$.
Now consider generally $\left(s-1\right)\ell\leq m\leq s\ell-1$ and $\left(t-1\right)\ell\leq n\leq t\ell-1$ with $n+m\leq\ell^{2}-1$. Then $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=ts$ and $b^{\ell}\left(n+m\right)=t+s$ if $n+m\geq\left(t+s-1\right)\ell$ and $b^{\ell}\left(n+m\right)=t+s-1$ if $n+m\leq\left(t+s-1\right)\ell-1$. Now we obtain $ts=t+s$ if and only if $\left(t-1\right)\left(s-1\right)=ts-t-s+1=1$ i. e. $t=2=s$ and ts=t+s1ts=t+s-1 if and only if $\left(t-1\right)\left(s-1\right)=ts-t-s+1=0$ i. e. $t=1$ or $s=1$. Similarly $ts<t+s$ if and only if $\left(t-1\right)\left(s-1\right)<1$ i. e. $t=1$ or $s=1$ and ts<t+s1ts<t+s-1 if and only if $\left(t-1\right)\left(s-1\right)<0$ i. e. never.
With PARI/GP we verified the conjecture for $\ell=6$ and $m,n\leq 3500$ and for $\ell=7$ and $m,n\leq 1200$.
Example 14.
Consider the following sequence consisting of $\alpha\left(n\right)=2^{n}+1$ for $n$ odd and $\alpha\left(n\right)=2^{n}+2$ for $n$ even. Then for $n$ odd we obtain
$\left(\alpha\left(n\right)\right)^{2}-\alpha\left(n-1\right)\alpha\left(n+1\right)$ $=$ $\left(2^{n}+1\right)^{2}-\left(2^{n-1}+2\right)\left(2^{n+1}+2\right)$
$=$ $2^{2n}+2^{n+1}+1-2^{2n}-2^{n+2}-2^{n}-4$
$=$ $-3\cdot 2^{n}-3<0$
which fulfills the log-convexity condition and for $n$ even we obtain
$\left(\alpha\left(n\right)\right)^{2}-\alpha\left(n-1\right)\alpha\left(n+1\right)$ $=$ $\left(2^{n}+2\right)^{2}-\left(2^{n-1}+1\right)\left(2^{n+1}+1\right)$
$=$ $2^{2n}+2^{n+2}+4-2^{2n}-2^{n-1}-2^{n+1}-1$
$=$ $3\cdot 2^{n-1}+3>0$
which fulfills the log-concavity condition. Therefore, the sequence is neither log-concave nor log-convex. Consider now for arbitrary $n$ $\left(\alpha\left(n\right)\right)^{n+1}\geq\left(2^{n}+1\right)^{n+1}=\left(2^{n}+1\right)^{n}\left(2^{n}+1\right)>\left(2^{n}+1\right)^{n}2^{n}=\left(2^{n+1}+2\right)^{n}\geq\left(\alpha\left(n+1\right)\right)^{n}$. For all $n$ therefore $\left(\alpha\left(n\right)\right)^{\frac{1}{n}}>\left(\alpha\left(n+1\right)\right)^{\frac{1}{n+1}}$ holds true. We can check that
$\alpha\left(n\right)\alpha\left(m\right)\geq\left(2^{n}+1\right)\left(2^{m}+1\right)=2^{n+m}+2^{n}+2^{m}+1>2^{n+m}+2\geq\alpha\left(n+m\right)$
for all $n,m$.
To summarize we have a sequence that is neither log-concave nor log-convex, satisfies the Bessenrodt–Ono inequality, and for which $\left(\alpha\left(n\right)\right)^{1/n}>\left(\alpha\left(n+1\right)\right)^{1/\left(n+1\right)}$.
Declarations and statements
Funding
This research received no external funding.
Competing interests
The authors declare that they have no competing interests.
Compliance with ethical standards
The authors declare that they comply with ethical standards
Data availability statement
No datasets were generated or analysed during the current study.
Author’s contribution statement
Both authors contributed in equal parts.